A sample of concentrated HCl solution with 88% by mass and density of 1.43 g cm⁻³. Calculate the molality of the solution


1
Expert's answer
2021-08-09T07:47:41-0400

HCl solution with 88% means

100g of Solution contains 88g of HCl

Moles of HCl = 88g/(36.45g/mol) = 2.41 mol

Density of solution = 1.43 gcm-3 = 1.43g/mL

100g solution of solution corresponds to

100g/(1.43g/mL) = 69.9 mL = 0.069 L

Molarity = moles / volume

Molarity = 2.41/0.069 = 34.92M